### v2

# Noise-based reward-modulated learning

Abstract
Biological neural systems efficiently learn from delayed rewards despite relying on noisy synaptic transmission and lacking centralized optimization mechanisms. In contrast, artificial neural networks trained with reinforcement learning typically rely on backpropagation (BP), which limits their use in resource-constrained systems or with non-differentiable components. While noise-based alternatives, like reward-modulated Hebbian learning (RMHL), provide a biologically grounded framework for credit assignment, they struggle with temporal delays and hierarchical processing —key challenges in real-world learning. In this work, we derive a novel noise-based learning rule to address these challenges. Drawing inspiration from biological neural circuits, our method uses reward prediction errors as its optimization target to generate increasingly advantageous behavior, and incorporates an eligibility trace to facilitate retrospective credit assignment. Its formulation relies on local information, aligning with biological constraints and enabling neuromorphic implementation. Experimental validation on reinforcement tasks (immediate and delayed rewards) shows our approach significantly outperforms RMHL and achieves performance comparable to BP, although with slower convergence due to its noise-driven updates. While tested on simple architectures, the results highlight the potential of noise-driven, brain-inspired learning for low-power adaptive systems, particularly in scenarios where energy efficiency and biological plausibility are a priority. These findings also offer mechanistic insights into how dopamine-like signals and synaptic stochasticity may jointly enable learning in biological networks, bridging computational models with neurobiological principles.
1 Introduction
The brain’s remarkable ability to learn and adapt to its environment is believed to arise from synaptic plasticity [4, 3, 37]. These local modifications allow neural circuits to acquire complex sensorimotor skills despite inherent challenges such as noisy neural communication [12], temporally delayed feedback/reward [23], and no centralized error-propagation mechanism to drive weight updates [48]. Understanding the mechanisms underlying this robust learning remains a central challenge in neuroscience and is increasingly inspiring novel approaches in artificial intelligence [11, 38].
When it comes to learning in artificial neural networks, backpropagation (BP) underpins modern deep learning [45]. However, its implementation in biological circuits is often seen as challenging [34]. BP requires precise, non-local error gradients, which contrasts with the locality of synaptic plasticity that depends on pre- and post-synaptic activity and neuromodulatory factors —akin to three-factor learning [21, 29]. Moreover, its deterministic nature is contrary to the inherent stochasticity of neural circuits [12, 51, 39], and its requirement for continuous, immediate error feedback contradicts biological learning, which often involves delayed feedback following a sequence of actions [33].
While some works have investigated how variants of backpropagation might be implemented in cortical microcircuits —employing predictive coding, feedback alignment, or dendritic error signals to carry nonlocal gradients [60, 59, 31]— others have proposed different plasticity rules inspired directly from neurobiology [46, 25, 63]. Among these approaches, noise‐based learning methods are especially promising. They leverage fluctuations in neural activity [19, 24, 61, 58, 16] or synaptic weights [13, 65, 7, 10], for estimating gradients. By injecting small random perturbations, these methods avoid the need for perfectly matched feedback pathways and rely only on information that biological synapses could plausibly access, aligning more closely with mechanisms observed in cortical neurons [50, 36].
Given the prevalence of noise in biological neural systems [12, 51, 39], using noise as a mechanism for learning synaptic weights is an area of growing interest, with reward-modulated Hebbian learning (RMHL) [32, 40] as a promising candidate. RMHL, which is a form of three-factor Hebbian learning [21, 29], offers a potentially powerful mechanism for credit assignment without explicit backpropagation of errors. Despite theoretical advances, few noise-based methods have been adapted to real-world tasks. Some attempts which have been made include application to control problems [6], for spatiotemporal pattern generation [28], and for training non-differentiable spiking neural networks training [15], and these particularly struggle with delayed reward regimes. This gap has limited the integration of noise-based approaches for effective credit assignment in temporally extended setups.
To bridge these gaps, we propose noise-based reward-modulated learning (NRL), a biologically-inspired, gradient-free learning method compatible with delayed feedback/rewards. NRL builds on established neuroscientific pillars. It employs eligibility traces [23, 26], which tag synapses based on recent pre/post activity and noise perturbations, enabling retrospective credit assignment over behavioral timescales. It also embraces reward prediction error (RPE), that is, dopamine-like signals [48, 47, 42] that modulate these traces to reinforce actions that yield unexpected positive rewards, driving learning toward increasingly more advantageous behaviors. We derive our learning rule from first principles, employing directional derivatives to compute a local gradient estimate at each synapse, rather than relying on explicit gradients. These directional derivatives are implemented through stochastic neurons, aligning with the inherently noisy nature of neural circuits and other physical systems. NRL estimates the directional derivative with two forward passes: a “noisy” pass with stochastic neurons and a “clean” pass without noise. In scenarios where a noiseless pass is infeasible, multiple noisy passes can be averaged to approximate the clean pass, maintaining performance.
We deliberately choose an abstract modelling framework so that our experiments isolate the learning rule’s behavior without introducing additional biological complexities. We employ rate-based neurons instead of spiking models, such that the dynamics reduce to smooth activation functions and continuous outputs. We also avoid distinguishing excitatory and inhibitory populations, allowing every unit to follow the same update equation. Finally, we treat potentiation and depression as two ends of a single weight-change continuum rather than separate processes, so that each synapse simply strengthens or weakens in proportion to its eligibility trace and reinforcement signal. In this way, we trade-off biological detail for more clear insight into capabilities and limitations of our proposed noise-based learning.
We validate NRL on a suite of reinforcement-learning benchmarks spanning immediate-reward and delayed-reward tasks, demonstrating significantly superior learning efficiency when compared to RMHL (particularly when rewards were delayed by many steps), and competitive final performance compared to gradient-based baselines. Moreover, unlike RMHL, which breaks down in deeper architectures, NRL scales to multilayer networks, though converging more slowly due to its intrinsic stochasticity. Finally, we explore how these findings shed light on biological learning and point toward energy-efficient implementations on neuromorphic hardware, as well as novel, alternative, robust learning schemes for machine learning.
2 Methods
2.1 Neural system-environment interaction and reward signaling
We model a neural system within a dynamic environment, grounded in the reinforcement learning framework [54]. In this framework, the environment’s state at time $t$ is denoted by $s_{t}$. The neural system perceives $s_{t}$ and responds by initiating actions, $a_{t}$, which influence the environment. Each action or sequence of actions results in feedback in the form of positive or negative rewards $r_{t}$. This iterative process of perception, action, and outcome forms a feedback loop that enables the system to learn and adapt to the changing environment, ultimately seeking to maximize the rewards it receives.
The system’s decision-making is guided by a policy $\pi$, representing the probability of choosing an action, $a_{t}$, in a given state $s_{t}$ at time $t$, i.e., $\pi(a_{t}\mid s_{t})$. This policy captures the system’s learned strategy for choosing actions that are likely to yield favorable outcomes. To determine an action, the system computes a probability distribution across the set of possible actions, modelled as a categorical distribution $P(a_{t}\mid s_{t})\equiv\pi(a_{t}\mid s_{t})$ with $\sum_{a\in A}\pi(a\mid s_{t})=1$, where $A$ denotes the set of all possible actions.
The rewards $r_{t}$, obtained from the environment, serve as crucial learning signals. Internally, the system maintains a prediction of the expected reward, denoted by $\bar{r}_{t}$. While this prediction can be calculated through various mechanisms, we model it here for simplicity as a running average of recent rewards:
$\bar{r}_{t+1}=\bar{r}_{t}+\lambda\left(r_{t+1}-\bar{r}_{t}\right)$ (1)
where $\lambda$ is a smoothing factor that governs the influence of past rewards on the current prediction. The mismatch between the actual reward and the predicted reward, known as the reward prediction error, is then computed as
$\delta_{t}=r_{t}-\bar{r}_{t}.$ (2)
This RPE signal is a central element of our model and draws a strong parallel to the phasic activity of dopamine neurons in the brain [48, 47, 42]. Dopamine neurons are known to exhibit increased activity for unexpected rewards and decreased activity for the unexpected absence of reward [17]. Although our running average is a simplified representation of how reward expectations are formed in biological systems, it effectively captures the fundamental dynamics of the RPE and its role as a learning signal. We will explore potential refinements of this reward prediction mechanism in the Discussion section.
2.2 Derivation of the learning rule
This section presents a complete derivation of the NRL update rule. We begin by establishing a gradient-based learning rule rooted in an optimization target that aims to maximize unexpected rewards from the environment by increasing the reward prediction error. Using RPE, rather than the direct reward, enables our system to adapt dynamically to changing environments by tracking how rewards deviate from expectations, improving long-term performance by continuously seeking rewards that exceed expectations, and maintaining robustness against variations in reward structures [54, 43]. Building on this foundation, we transition to a directional derivative framework, where stochastic noise within the network approximates gradients. This eliminates the need for explicit backpropagation and feedback phases, enabling learning through forward passes alone. Finally, we extend our noise-based approach to handle scenarios with delayed rewards, a hallmark of real-world problems. This extension enables NRL to adapt to environments where feedback is only obtained after a series of actions.
2.2.1 Gradient-based learning rule
The system is modelled as a multi-layered network of interconnected units, analogous to populations of neurons, which processes input representing sensory observations and produces a probability distribution over possible actions. For simplicity, we refer to these populations as ‘layers,’ where each layer performs a transformation of its input in two stages: first, a linear transformation, followed by a non-linear transformation. This can be expressed mathematically as
$x^{l}_{t}=f(h^{l}_{t})=f(W^{l}_{t}x^{l-1}_{t})$ (3)
where $x^{l}_{t}$ represents the output activity of layer $l$, $h^{l}_{t}$ represents the layer pre-activation (the combined input to such layer), $W^{l}_{t}$ represents the matrix of synaptic weights connecting layer l1l-1 to layer $l$, at time $t$, and $f(\cdot)$ is a non-linear activation function, which introduces crucial non-linearities into the network’s computations
In this setup, the goal is to adjust the synaptic weights, $\{W^{1},\ldots,W^{L}\}$, across the $L$ layers of the network to maximize the rewards obtained by the system. Our derivation begins by introducing a general parameter $\theta$, which will later be mapped to the specific neural network parameters $W$. The primary learning objective is to maximize the RPE, $\delta_{t}$, at any particular moment in time, $t$, effectively driving the system to seek out actions that lead to higher-than-predicted rewards. The reward prediction is modeled as a running average of recent rewards and, as such, the learning rule seeks to outperform previously received rewards. Note that for a fixed reward prediction, this is equivalent to maximizing the reward itself. We express this objective in terms of a parameterized policy at time $t$ as
$J(\theta_{t})=\mathbb{E}_{\pi_{\theta_{t}}}\left[\delta_{t}\right]\,.$ (4)
To optimize $J(\theta_{t})$, we can incrementally update the weights using the gradients with respect to $\theta_{t}$ as
$\theta_{t+1}\leftarrow\theta_{t}+\eta\nabla J(\theta_{t})$ (5)
where $\eta$ is the learning rate, controlling the size of each update step.
Applying the policy gradient theorem [55], which utilizes the likelihood-ratio method, we express the gradient of the objective as
$\nabla J(\theta_{t})=\mathbb{E}_{\pi_{\theta_{t}}}\left[\nabla\log\pi_{\theta_{t}}\left(a_{t}\mid s_{t}\right)\delta_{t}\right]\,.$
For empirical estimation, in the case of a single sample, we approximate the gradient as $\nabla J(\theta_{t})\approx\nabla\log\pi_{\theta_{t}}\left(a_{t}\mid s_{t}\right)\delta_{t}$. Using this approximation, we define the parameter updates as
$\theta_{t+1}\leftarrow\theta_{t}+\eta\;\nabla\log\pi_{\theta_{t}}\left(a_{t}\mid s_{t}\right)\;\delta_{t}\,.$ (6)
Equation (6) resembles the REINFORCE update rule [61] but differs by using the reward prediction error, $\delta_{t}$, as the learning signal instead of cumulative rewards over full trajectories. This RPE-based approach leverages immediate feedback from rewards as they are obtained rather than requiring a full trial completion to estimate the policy gradient. It shares conceptual similarities with actor-critic methods in reinforcement learning [1], where the policy is adjusted using a temporal difference error. However, we approximate future rewards with a running average of past rewards instead of a critic network, a less powerful but simpler implementation which maintains adaptive feedback.
2.2.2 Noise-based learning rule
The learning rule derived in the previous section still relies on gradient descent. To avoid using backpropagation to compute the gradients, we propose a noise-based alternative that extends Equation (6). This approach leverages gradient approximation via directional derivatives, enabling a theoretically rigorous derivation of noise-driven learning.
A directional derivative quantifies the rate of change of a function in a specified direction. In our neural system, we implement this concept by introducing random noise into parameters. By comparing the network’s parameters with and without this noise, we obtain an estimate of the gradient direction. To formalize this, we define $g(\theta_{t})=\log\pi_{\theta_{t}}(a_{t}\mid s_{t})$ and express the gradient term in terms of directional derivatives using the theorems in Appendix A as
$\nabla g(\theta_{t})$ $=\nabla\log\pi_{\theta_{t}}(a_{t}\mid s_{t})=n\mathbb{E}\left[\bar{\epsilon}_{t}\nabla_{\bar{\epsilon}_{t}}g(\theta_{t})\right]$ (7)
[…]
We may expand the above to approximate the gradient via a finite-difference and by sampling under an empirical distribution
[…]
where $K$ denotes the number of samples and $\tilde{\theta}_{t}^{(i)}=\theta_{t}+\epsilon^{(i)}_{t}$ the noise-perturbed parameters (see Appendix A). In practice, we consider $K=1$, analogous to single-sample updates in stochastic gradient descent. Here, $g\bigl{(}\tilde{\theta}^{(i)}_{t}\bigr{)}=\log\pi_{\tilde{\theta}_{t}}(a_{t}\mid s_{t})$ and $g\left(\theta_{t}\right)=\log\pi_{\theta_{t}}(a_{t}\mid s_{t})$ are the log of the noise-perturbed and noise-free output, respectively. Thus, we define $\rho_{t}=\log\pi_{\tilde{\theta}_{t}}(a_{t}\mid s_{t})-\log\pi_{\theta_{t}}(a_{t}\mid s_{t})$, which captures the impact of the noise on the policy. Putting this together for the $K=1$ case, we obtain our noise-based learning rule
$\theta_{t+1}\leftarrow\theta_{t}+\eta\delta_{t}\hat{\epsilon}_{t}\rho_{t}$
[…]
2.2.3 Learning with node level noise in neural network
The noise-based learning rule derived so far is formulated for a general parameter $\theta_{t}$. Now, we apply this framework to the specific context of a neural network, where $\theta_{t}$ corresponds to the synaptic weights $W^{l}_{t}$ of layer $l$ at time $t$.
Instead of directly perturbing the weights of the network $W^{l}_{t}$, we propose extending the learning rule by introducing noise directly into the neurons (nodes) of each layer. This strategy is advantageous because it can lead to reduced variance in gradient estimation, as perturbations occur in a lower-dimensional space (the neural activations) compared to the full weight space. This approach shares similarities with node perturbation (NP) [19, 24, 61, 58, 16], which has shown benefits in reducing gradient variance and offering more localized updates.
Given a layer transformation as in Equation (3), adding noise at the neuron level is represented as
$\tilde{x}^{l}_{t}=f\left(\tilde{h}^{l}_{t}+\xi^{l}_{t}\right)=f\left(W^{l}_{t}\tilde{x}^{l-1\top}_{t}+\xi^{l}_{t}\right)$ (9)
where $\xi^{l}_{t}\sim\mathcal{N}(0,\sigma^{2}I_{m^{l}})$ is the injected noise at time $t$, $\sigma^{2}$ is some arbitrarily small noise scale, and $m^{l}$ is the number of neurons in layer $l$. We use the notation $\tilde{x}^{l}_{t}$ and $\tilde{h}^{l}_{t}$ to denote perturbed inputs and pre-activations, respectively, which may also result from perturbations of previous network layers on which layer $l$ depends.
To formulate the learning rule for specific layer parameters $W^{l}_{t}\in\{W^{1}_{t},\ldots,W^{L}_{t}\}$, we re-express the gradient in terms of the pre-activations instead of the parameters themselves. Here, the general parameter vector $\theta_{t}=\{W^{1}_{t},\ldots,W^{L}_{t}\}$ corresponds to the full set of all weight matrices for all layers at time $t$. To do so, we first rewrite the gradient term as
$\nabla g\left(W^{l}_{t}\right)=\nabla\log\pi_{\theta_{t}}\left(a_{t}\mid s_{t}\right)$ (10)
and apply the chain rule to break down the gradient with respect to $W^{l}_{t}$ as
$\nabla g\left(W^{l}_{t}\right)=\nabla g\left(h^{l}_{t}\right)\frac{\partial h^{l}_{t}}{\partial W^{l}_{t}}=\nabla g\left(h^{l}_{t}\right){{x}^{l-1\top}_{t}}.$ (11)
Notably, we may once again carry out the conversion from this gradient estimation step to a set of directional derivatives such that
[…]
where $i$ indexes over a set of repeated samples of the noise term, $\xi$.
Again, as above, by reducing this to the single sample ($n=1$) case, we can define a layer-specific weight update rule
$W^{l}_{t+1}\leftarrow W^{l}_{t}+\eta\delta_{t}\bar{\xi}^{l}_{t}{\rho}_{t}{\tilde{x}^{l-1\top}_{t}}$ (13)
[…]
2.2.4 Learning from sparse rewards
Our derivations so far assume that synaptic updates happen at every time step, implying a continuous stream of rewards and learning signals. In reality, however, rewards are often sparse, arriving only after a sequence of actions or upon reaching specific milestones. To handle this, we’ll modify our learning rule to account for rewards received at arbitrary times.
Let’s denote these discrete reward times as $\tau_{m}\in\{\tau_{0},\tau_{1},...\tau_{M}\}$. At these moments, we compute the RPE, $\delta_{\tau_{m}}$, as
$\delta_{\tau_{m}}=r_{\tau_{m}}-\bar{r}_{\tau_{m}}.$ (14)
Here, $r_{\tau_{m}}$ is the reward received at time $\tau_{m}$, and $\bar{r}_{\tau_{m}}$ is the reward prediction, which is updated as the running average of recent rewards.
$\bar{r}_{\tau_{m}}=\bar{r}_{\tau_{m-1}}+\lambda\left(r_{\tau_{m}}-\bar{r}_{\tau_{m-1}}\right)$ (15)
While synaptic updates only occur at these specific reward times $\tau_{m}$, the learning rule continuously accumulates information between rewards. This information is integrated over time, allowing us to update weights based on what’s happened since the last reward. This leads to our modified learning rule
$W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\delta_{\tau_{m}}\;e_{\tau_{m}}$ (16)
where
$e_{\tau_{m}}=\sum_{t=\tau_{m-1}}^{\tau_{m}}\bar{\xi}^{l}_{t}{\rho}_{t}{\tilde{x}^{l-1\top}_{t}}$ (17)
is an eligibility trace that acts as a mechanism to connect past actions with future rewards [23]. Eligibility traces capture neural activity and other local variables over time, signaling potential synaptic changes. Upon receiving a reward, these traces are modulated by the reinforcement signal, resulting in synaptic updates. Some models view eligibility traces as decaying cumulative activity [26, 54]), while others treat them as a full activity history [40], which aligns with our formulation. The eligibility trace efficiently tracks neural information between rewards, facilitating the assignment of credit to past actions.
Thus, our final learning rule consists of two primary components: (i) an eligibility trace, defined in Equation (17), which accumulates local information over time at each time step, and (ii) a synaptic update, defined in Equation (16), triggered upon reward receipt, which modulates the eligibility trace to adjust the synaptic weights. This learning rule constitutes the core of our proposed noise-based reward-modulated Learning (NRL) and is used for all experiments presented in this paper.
2.3 Neural network architecture
The learning and decision-making capabilities of the system are modelled using a feedforward network with $L$ layers, representing interconnected populations of neurons. The transformation performed by each hidden layer at time $t$ can be expressed as
$\tilde{x}^{l}_{t}$ $=f\left(W^{l}_{t}\tilde{x}^{l-1\top}_{t}+\xi^{l}_{t}\right)$ (18)
and the readout in the output layer is given by
$\tilde{y}_{t}$ $=s\left(W^{l}_{t}\tilde{x}^{l-1\top}_{t}+\xi^{l}_{t}\right)$ (19)
where $x^{l}_{t}$ represents the activity of units in layer $l$, $W^{l}_{t}$ is the matrix of synaptic weights connecting layer l1l-1 to layer $l$, and $\xi^{l}_{t}\sim\mathcal{N}(0,\sigma^{2}I_{m^{l}})$ represents random noise introduced into layer $l$, with $m^{l}$ being the number of units in such layer. This noisy propagation of activity is termed a “noisy pass.” The function $f(\cdot)$ is a non-linear activation function, which we implement as the LeakyReLU function such that $f(x)=x$ if $x\geq 0$ and $f(x)=\alpha x$ with $0<\alpha\ll 1$. The function $s(\cdot)$ represents a softmax transformation ${s}_{i}(x)=\ifrac{e^{x_{i}}}{\sum_{j=1}^{m^{l}}e^{x_{j}}}$, which transforms the network’s output into a probability distribution over possible actions. We also define a “clean pass,” representing the network’s activity in the absence of noise, by $x^{l}_{t}=f\left(W^{l}_{t}x^{l-1\top}_{t}\right)$ and $y_{t}=s\left(W^{l}_{t}x^{l-1\top}_{t}\right)$.
2.4 Experimental validation
We validate our approach across a range of simulated environments, comparing it against established baseline learning methods. Each environment presents a problem with discrete episodes, where the system’s learning is guided solely by positive or negative rewards. We investigate both scenarios with immediate reward and those involving delayed reward. In the delayed reward scenarios, a single reward is provided after a sequence of actions, challenging our system to assign credit retrospectively to the actions that contributed to the outcome. For the instantaneous reward setting, we use the Reaching problem [22], while the delayed reward setting uses the Cartpole [1] and Acrobot [53] problems. We use implementations given by the libraries OpenAI Gym [5] and NeuroGym [41] for the different environments.
We compare NRL to two baselines: an exact-gradient version of NRL, which serves as an ”optimal performance” benchmark, and an RMHL approach.
The first baseline, which we refer to as the exact-gradient method (BP baseline), is similar to an actor-only variant of actor-critic methods, relying on a running average of past rewards as the prediction error. For this baseline, we compute the exact policy gradients for the entire episode trajectory using backpropagation through time (BPTT) [57]. Given sparse rewards, the weights at each layer are updated at the end of each episode according to
$W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\left(\nabla_{W^{l}_{\tau_{m-1}}}\log\pi_{\theta_{\tau_{m-1}}}(a_{t}\mid s_{t})\right)\delta_{\tau_{m}}$ (20)
The second baseline, motivated by the noise-based nature of NRL, is the RMHL rule from [32] with an explicit-noise approximation – the original version where noise is inferred from neural activities proved too unstable for the problems considered here. For delayed rewards, we adapt it similarly to [40], updating eligibility traces as
$e_{\tau_{m}}=\sum_{t=\tau_{m-1}}^{\tau_{m}}\xi^{l}_{t}{\tilde{x}^{l-1\top}_{t}}\,.$ (21)
Similar to NRL’s noise-based learning rule defined in Equation (16), the synaptic weights for RMHL are then updated by modulating this eligibility trace with the reward prediction error.
In Table 1, we present the complete three learning rules (NRL, BPTT, and RMHL) in a single table for visualization and quick comparison. All updates are written in a per-layer form using $W^{l}$ and assume a delayed reward received at time $\tau_{m}$.
Table 1: Comparison of the learning rules.
Rule Update
NRL $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\delta_{\tau_{m}}\!\sum_{t=\tau_{m-1}}^{\tau_{m}}\bar{\xi}^{\,l}_{t}\,\rho_{t}\,\tilde{x}^{\,l-1\top}_{t}$
BPTT $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\left(\nabla_{W^{l}_{\tau_{m-1}}}\log\pi_{\theta_{\tau_{m-1}}}(a_{t}\mid s_{t})\right)\delta_{\tau_{m}}$
RMHL $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\delta_{\tau_{m}}\!\sum_{t=\tau_{m-1}}^{\tau_{m}}\xi^{\,l}_{t}\,\tilde{x}^{\,l-1\top}_{t}$
In environments with high reward variability, like Cartpole and Acrobot, we stabilize synaptic updates by scaling the RPE by dividing by $r_{\tau_{m}}$. Hyperparameter values and training details are provided in Appendix B. All experiments are implemented in Python using PyTorch [44]. Our models and scripts are available for reproducibility at https://github.com/jesusgf96/noise-based-reward-modulated-learning.
3 Results
3.1 Solving instantaneous and delayed reward problems
To ensure comparability with RMHL methods, which typically utilize single-hidden-layer networks, we first conduct experiments with a one-hidden-layer neural network for BP, NRL, and RMHL. Training details are provided in Appendix B. In Section titled “Scalability to multi-layer networks”, we extend the comparison to deeper networks with multiple hidden layers.
First, the Reaching problem [22], visualized in Fig. 1C, is an immediate reward problem that requires the system to reach and maintain a position at a target on a 1D ring by moving left, right, or remaining stationary. At each step, the system receives information about both the target’s position and its own, and a reward is provided based on proximity to the target over fixed-duration trials. Average performance across-trial and final performance (mean of the last 50 trials) are shown in Fig 1A and Fig 1B, respectively.
Figure 1: Performance on benchmarks. A, B, C: Reaching problem. D, E, F: Acrobot problem. G, H, I: Cartpole problem. Left panels: Performance across trials averaged over 5 runs. Centre panels: Final performance (mean of the last 50 trials), averaged over 5 runs. Right panels: Problem visualization.
Second, the Acrobot problem [53], visualized in Fig 1F, involves delayed reward and requires controlling a two-link robotic arm to reach a target height. At each time step, the system receives information about the angles and angular velocities of the two links and chooses one action: clockwise torque, counterclockwise torque, or no torque. Rewards are given based on the speed of completion, with a maximum time allowed. Average performance across trials and final performance (mean of the last 50 trials) are shown in Fig 1D and Fig 1E, respectively.
The third and most challenging problem, the Cartpole problem [1], visualized in Fig 1I, is a delayed reward problem where the system must balance a pole on a cart by moving left or right. At each time step, the system receives information about the cart’s position and velocity, along with the pole’s angle and angular velocity, and responds accordingly. Performance is measured by the time the pole remains balanced, with a maximum time allowed. Average performance across trials and final performance (mean of the last 50 trials) are shown in Fig 1G and Fig 1H, respectively. For all three tasks, NRL reaches the performance level of BP, whereas RMHL fails to do so.
3.2 Scalability to multi-layer networks
Here, we demonstrate that NRL can effectively assign credit in neural networks with multiple hidden layers; a challenging scenario where most biologically plausible algorithms struggle. We use both the Acrobot and the Cartpole tasks for this purpose, as they present a delayed reward problem, making them a more realistic test for credit assignment in reinforcement learning. In this comparison, we include the same baselines, BP and RMHL. Fig 2 displays the results for the Acrobot task and Fig 3 displays the results for the Cartpole tasks. The left panels of these two figures show performance across trials, averaged over 5 runs. In contrast, the right panels display the final performance (mean of the last 50 trials), also averaged over 5 runs. The results for neural networks with two hidden layers are shown in Fig 2A and Fig 2B, and 3A and B, and those for networks with three hidden layers are presented in Fig 2C and Fig 2D, and Fig 3C and Fig 3D.
Figure 2: Performance on deeper networks for the Acrobot. A, B: 2-hidden layer networks. C, D: 3-hidden layer networks. Left panels: Performance across trials averaged over 5 runs. Right panels: Final performance (mean of the last 50 trials), averaged over 5 runs.
Figure 3: Performance on deeper networks for the Cartpole. A, B: 2-hidden layer networks. C, D: 3-hidden layer networks. Left panels: Performance across trials averaged over 5 runs. Right panels: Final performance (mean of the last 50 trials), averaged over 5 runs.
Our results indicate that NRL successfully learns to solve the tasks, achieving performance comparable to BP. In contrast, RMHL struggles with credit assignment in deeper networks. However, NRL requires more trials to converge as the network depth increases – an expected outcome due to the stochastic nature of the updates [24]. A similar trend is observed with BP, though to a lesser extent, as its gradient-based updates inherently provide more directed adjustments.
3.3 Using only noisy passes
In NRL, $\rho_{t}$ is calculated as the difference in the network’s output between the clean and noisy passes. However, a completely noiseless network is not feasible in more biophysically realistic scenarios. Instead, the clean network’s output can be approximated by averaging the outputs from multiple noisy passes, as the injected noise averages to zero $\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\xi_{t,i}^{l}=0$. This assumes that the network dynamics are faster than the environment dynamics, allowing the network to perform multiple forward passes before the environment changes.
To evaluate the accuracy of this approximation, we employ a single hidden-layer network and the Acrobot problem. The difference between the clean pass output and the averaged noisy passes output was calculated for each timestep and averaged over 500 timesteps, as shown in Fig 4A. Furthermore, Fig 4B illustrates the performance on the Acrobot problem using only noisy forward passes, starting with a minimum of 2 noisy passes. We also extended this evaluation to 10 noisy passes to explore the impact of increasing the number of passes, showing slightly faster initial convergence.
Figure 4: Learning using only noisy passes for the Acrobot. A: Clean pass approximation error. Each data point is computed with an absolute error, averaged over 500 timesteps using the Acrobot problem. B: Acrobot problem. Performance across trials, averaged over 5 runs, with mean, minimum, and maximum values displayed.
4 Discussion
In this work, we propose NRL, a noise-driven synaptic plasticity rule that bridges principles of reinforcement learning and optimization with biologically-inspired Hebbian updates. By leveraging stochastic neural activity, our method produces local synaptic updates modulated by a global reinforcement signal. It offers an efficient approach for neural adaptation, particularly suited to neuromorphic systems.
We employ a top-down methodology, starting by deriving our learning rule from the mathematical principles of gradient-based optimization. We then approximate the gradients via directional derivatives and incorporate bottom-up constraints inspired by biological systems. Specifically, the gradient-approximation process is implemented through stochastic neural activity (aligning with the inherent noise in the brain [12, 51, 39]), and reinforcement learning mechanisms are realized through reward prediction errors (emulating dopaminergic reward signals [48, 47]), eligibility traces, and local synaptic plasticity. These elements enable learning under uncertainty [18]. This approach results in a learning rule that is both theoretically grounded and potentially realizable in neural circuits.
Our results demonstrate the learning effectiveness of NRL, solving both immediate and delayed feedback/reward problems in simulated environments, achieving performance comparable to baselines optimized employing backpropagation. Additionally, it shows a significant improvement over RMHL [32, 40], a modern, well-known approach that employs noise as the core of its synaptic plasticity. Our scalability results further demonstrate that NRL effectively assigns credit even in multilayer networks, a challenge for many biologically plausible algorithms, indicating its potential for scaling. This is a critical difference with RMHL, which struggles to assign credit in multilayer networks evidenced by its minimal performance improvements across trials. Nevertheless, as the depth of the network increases, NRL shows slower convergence due to its noise-driven updates.
NRL shares similarities with node perturbation methods [19, 24, 61, 58, 16], which inject noise into neurons to adjust weights based on state and performance fluctuations. However, unlike node perturbation – which requires constant feedback or reward for learning – NRL is well-suited for delayed reward setups. By leveraging eligibility traces [23, 26], it can effectively assign credit to past neural states even when actions and outcomes are separated in time.
Gradient approximation, a cornerstone of our learning rule’s derivation, is also crucial in mathematical optimization and machine learning, enabling learning when exact gradients are computationally expensive or unavailable. Techniques such as forward gradient methods [2] and zeroth-order (ZO) optimization [35, 8] have emerged, utilizing stochastic perturbations and function value comparisons to estimate gradients. Recent advances, such as employing local auxiliary networks for informed gradient guesses [20], have significantly improved the alignment of these approximations with exact gradients, reducing the performance gap with backpropagation. These variance-reduction strategies could similarly enhance our approach by improving the quality of the updates and scalability to more complex architectures. Furthermore, compared to traditional gradient-based methods like backpropagation, gradient approximation methods offer a key advantage in computationally constrained scenarios, or when using non-differentiable networks, such as spiking neural networks [62, 56].
Additionally, the local-update structure of NRL provides compatibility with neuromorphic substrates [49], where efficient, low-power computations are essential. Neuromorphic systems can also inherently simulate noisy neural activity, enabling energy-efficient learning in applications where real-time adaptation is key. The inherent noise in NRL is also advantageous for reinforcement learning setups, as it naturally facilitates exploration-exploitation balance [27, 30]. Injected noise perturbs neural states, promoting exploratory deviations from expected values and thus introducing variations in the agent’s policy.
Our work has limitations that suggest future research directions. First, its dependence on multiple forward passes per synaptic update reduces biological plausibility, even though these passes could be performed in parallel. Future work should explore more efficient ways to estimate noise impacts, potentially eliminating the need for multiple passes. Second, our modeling approach deliberately involves a high degree of abstraction: we use rate-based neurons instead of spiking models, do not distinguish between excitatory and inhibitory populations, and treat synaptic potentiation and depression as a single continuum. While this provides clear insight into the capabilities and limitations of our approach, more biologically detailed models could offer further insights. Third, our reward prediction error calculation is a simplification. While sufficient for demonstrating the learning mechanism, the simplicity of using a running average may limit its effectiveness in scenarios where navigating through transient low-reward regions is necessary to achieve higher long-term rewards. In such cases, the model may underestimate the value of actions that lead to delayed benefits, potentially resulting in suboptimal updates. More sophisticated approaches, such as temporal difference learning [52, 54], or learned prediction through a critic network, could provide a richer and more temporally sensitive estimate of expected rewards. These strategies, which align more closely with biological dopaminergic signaling and can be implemented within actor-critic architectures, represent promising directions for improving learning performance in future work. Finally, while backpropagation may offer faster convergence in unconstrained settings, recent studies [9, 14], including ours, indicate that this gap can be narrowed for specific networks and tasks.
Our findings highlight the potential of unifying optimization methods and neurobiological principles. By drawing inspiration from both domains, we have developed a learning rule that not only offers a biologically plausible alternative to backpropagation but also provides insights into how learning, particularly synaptic plasticity, might occur in biological neural systems. Conversely, the constraints of biological realism can guide the development of more efficient and robust artificial learning algorithms, particularly suited for neuromorphic hardware and real-time applications. This synergistic approach promises to advance our understanding of intelligence in both biological and artificial systems [64, 38], driving progress in neuroscience and machine learning alike.
Acknowledgments
This publication is part of the DBI2 project (024.005.022, Gravitation), which is financed by the Dutch Ministry of Education (OCW) via the Dutch Research Council (NWO).
Appendix A Gradient approximation using directional derivatives
Consider a function $g(\theta)$, where $\theta=(W_{1},\ldots,W_{L})$ are its parameters. The gradient of this function is defined as follows:
Definition A.1.
The gradient $\nabla g$ is a vector indicating the direction of the steepest ascent of the function $g$, with components as partial derivatives of $g(\theta)$:
$\nabla g=\left[\frac{\partial g}{\partial\theta}\right]^{\top}=\left[\frac{\partial g}{\partial W_{1}},\ldots,\frac{\partial g}{\partial W_{L}}\right]^{\top}\,.$ (22)
[…]
Definition A.2.
The directional derivative of $g(\theta)$ along a unit vector $v=(v_{1},\ldots,v_{n})$ is defined by the limit
$\nabla_{v}{g}(\theta)=\lim_{h\to 0}{\frac{g(\theta+hv)-g(\theta)}{h}}\,$, (23)
where $h$ is a small step size.
For a sufficiently small $h$ we numerically measure the directional derivative as
[…]
This directional derivative can also be measured as a projection of $\nabla g$ in the direction $v$, following the relation:
$\nabla_{v}{g}(\theta)=v\cdot\nabla{g}(\theta).$ (25)
We can now formally demonstrate how gradients can be approximated using directional derivatives.
Theorem A.3.
Let $\epsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$ (probability distribution $p(\epsilon)$) where $n$ is the number of dimensions in $\theta$. Exact gradients can be written in terms of directional derivatives using expectations
$\nabla g(\theta)=n\mathbb{E}_{p(\epsilon)}\left[v^{\top}\nabla_{v}g(\theta)\right]\,.$ (26)
Proof.
$\nabla g(\theta)$ $=n\mathbb{E}_{p(\epsilon)}\left[v^{\top}\nabla_{v}g(\theta)\right]\,$
$=n\mathbb{E}_{p(\epsilon)}\left[vv^{\top}\nabla g(\theta)\right]\,$
$=n\mathbb{E}_{p(\epsilon)}\left[vv^{\top}\right]\nabla g(\theta)\,$
[…]
$=n\frac{1}{n}I_{n}\nabla g(\theta)=\nabla g(\theta)\,$
[…]
Now, consider gradient descent in the direction of the gradient $\nabla g(\theta)$ as
$\theta_{t+1}=\theta_{t}+\alpha\nabla g(\theta_{t}).$ (27)
This update can be reformulated using directional derivatives.
Theorem A.4.
Let $\eta=\frac{\alpha n}{hK}$ and $\epsilon^{(i)}\sim\mathcal{N}(0,\sigma^{2}I_{n})$. Gradient descent is equivalent to the update rule
[…]
in the limit when $h\to 0$ and $K\rightarrow\infty$.
Proof.
$\nabla g(\theta)$ $=n\mathbb{E}_{p(\epsilon)}\left[v\nabla_{v}g(\theta)^{\top}\right]\,$
[…]
Now, substituting the expectation with a sampling under some empirical distribution,
[…]
Finally, defining gradient descent on parameters $\theta$ as $\theta_{t+1}=\theta_{t}+\alpha\nabla g(\theta)$,
$\theta_{t+1}$ $=\theta_{t}+\alpha\nabla g(\theta)$
[…]
In practice, small perturbations, controlled by $h$, and limited numbers of noise samples are sufficient to approximate the gradient. For $K=1$, analogous to single-sample updates in stochastic gradient descent:
[…]
with $\rho=\left[g(\theta+h\epsilon)-g(\theta)\right]^{\top}$.
Appendix B Model hyperparameters and training details
All hyperparameters were carefully tuned per method and problem to ensure fair comparisons across methods. The number of episodes was chosen to illustrate the convergence of our method relative to baseline methods, with 8000, 20000, and 1000 episodes used for the Acrobot, Cartpole, and Reaching problems, respectively. In all our experiments, we set the smoothing factor for reward estimation to $\lambda=0.66$, striking a balance between recent values (for quick adaptability) and a longer history (for robustness against rapid reward fluctuations). Each neural network consisted of an input, hidden, and output layer. Input units corresponded to environment observation elements, and output units to possible actions. Specifically, Acrobot, Cartpole, and Reaching used 6, 4, and 32 input units and 3, 2, and 2 output units, respectively, with hidden layer sizes of 64, 64, and 128. Table 2 summarizes the learning rate $\eta$ and noise standard deviation $\sigma$ for each method and problem. Higher values of these parameters could lead to unstable training, while lower values may result in slower learning.
Table 2: Learning rate $\eta$ and noise standard deviation $\sigma$ for the different learning algorithms across problems. Dashes indicate that the parameter is not used.
BP Ours RMHL
Acrobot $\eta$ 5e-3 5e-2 5e-2
$\sigma$ - 1e-3 1e-3
Cartpole $\eta$ 5e-3 5e-2 1e-2
$\sigma$ - 1e-3 1e-1
Reaching $\eta$ 1e-2 1e-2 1e-1
$\sigma$ - 1e-3 1e-1